\begin{thebibliography}{10}



\bibitem{BrowderPetryshyn1967inverseStronglyMonotone}
F.E. Browder and W.V. Petryshyn.
\newblock Construction of fixed points of nonlinear mappings in Hilbert space.
\newblock {\em J. Math. Anal. Appl.}, 20:197--228, 2005.
%\newblock MR0217658.

\bibitem{[1]CaiBu2011strongConvergence}
G.~Cai and S.~Bu.
\newblock Strong convergence theorems based on a new modified extragradient
  method for variational inequality problems and fixed point problems in Banach
  spaces.
\newblock {\em Computers and Mathematics with Applications}, 62:2567--2579,
  2011.

\bibitem{[2]Chantaragsi2010coerciveMappings}
W.~Chantarangs, C.~Jaiboon, and P.~Kumam.
\newblock A viscosity hybrid steepest descent method for generalized mixed
  equilibrium problems and variational inequalities for relaxed cocoercive
  mapping in Hilbert spaces.
\newblock {\em Abstract and Applied Analysis}, Art. ID 390972:39 pp., 2010.
%\newblock MR2669084.

\bibitem{[3]Chen2007nonexpansiveMappings}
J.~Chen, L.~Zhang, and T.~Fan.
\newblock Viscosity approximation methods for nonexpansive mappings and
  monotone mappings.
\newblock {\em J. Math. Anal. Appl.}, 334(2):1450--1461, 2007.
%\newblock MR2338673.

\bibitem{[4]IidukaTakahashi2005strongConvergenceTheorems}
H.~Iiduka and W.~Takahashi.
\newblock Strong convergence theorems for nonexpansivemappings and
  inverse-strongly monotone mappings.
\newblock {\em Nonlinear Anal.}, 61(3):341--350, 2005.
%\newblock MR2123081.



\bibitem{[6]Jaiboon2010cocoerciveMappings}
C.~Jaiboon and P.~Kumam.
\newblock Strong convergence for generalized equilibrium problems, fixed point
  problems and relaxed cocoercive variational inequalities.
\newblock {\em J. Inequal. Appl.}, Art. ID 728028:43 pp., 2010.
%\newblock MR2600206.

\bibitem{[7]Jitpeera2011fixedPointProblems}
T.~Jitpeera and P.~Kuma.
\newblock A new hybrid algorithm for a system of mixed equilibrium problems,
  fixed point problems for nonexpansive semigroup, and variational inclusion
  problem.
\newblock {\em Fixed Point Theory Appl.}, Art. ID 217407:27 pp., 2011.
%\newblock MR2764772.

\bibitem{[9]Jitpeera2011hybridAlgorithms}
T.~Jitpeera and P.~Kumam.
\newblock Hybrid algorithms for minimization problems over the solutions of
  generalized mixed equilibrium and variational inclusion problems.
\newblock {\em Math. Probl. Eng.}, Art. ID 648617:25 pp., 2011.
%\newblock MR2826902.

\bibitem{[8]Jitpeera2011equilibriumProblems}
T.~Jitpeera and P.~Kumam.
\newblock A new hybrid algorithm for a system of equilibrium problems and
  variational inclusion.
\newblock {\em Ann. Univ. Ferrara Sez. VII Sci. Mat.}, 57(1):89--108, 2011.
%\newblock MR2821372.

\bibitem{[10]Jitpeera2011shrinkingProjection}
T.~Jitpeera and P.~Kumam.
\newblock The shrinking projection method for common solutions of generalized
  mixed equilibrium problems and fixed point problems for strictly
  pseudocontractive mappings.
\newblock {\em J. Inequal. Appl.}, Art. ID 840319:25 pp., 2011.
%\newblock MR2781420.

\bibitem{[11]Jung2011compositeIterativeMethod}
J.S. Jung.
\newblock A general composite iterative method for generalized mixed
  equilibrium problems, variational inequality problems and optimization
  problems.
\newblock {\em J. Inequal. Appl.}, 51:23 pp., 2011.
%\newblock MR2837906.

\bibitem{[12]Kangtunyakarn2010iterativeMethods}
A.~Kangtunyakarn.
\newblock Iterative methods for finding common solution of generalized
  equilibrium problems and variational inequality problems and fixed point
  problems of a finite family of nonexpansive mappings.
\newblock {\em Fixed Point Theory Appl.}, Art. ID 836714:29 pp., 2010.
%\newblock MR2747251.

\bibitem{[14]Kazmi2018hybridIterativeMethod}
K.R. Kazmi, Rehan Ali, and Mohd Furkan.
\newblock Hybrid iterative method for split monotone variational inclusion
  problem and hierarchical fixed point problem for a finite family of
  nonexpansive mappings.
\newblock {\em Numer. Algorithms}, 79(2):499--527, 2018.
%\newblock MR3857923.

\bibitem{[13]Kazmi2012extragradientMethod}
K.R. Kazmi and S.H. Rizvi.
\newblock A hybrid extragradient method for approximating the common solutions
  of a variational inequality, a system of variational inequalities, a mixed
  equilibrium problem and a fixed point problem.
\newblock {\em Appl. Math. Comput.}, 218(9):5439--5452, 2012.
%\newblock MR2870065.

\bibitem{[15]Kumam2009newHybridIterativeMethod}
P.~Kumam and C.~Jaiboon.
\newblock A new hybrid iterative method for mixed equilibrium problems and
  variational inequality problem for relaxed cocoercive mappings with
  application to optimization problems.
\newblock {\em Nonlinear Anal. Hybrid Syst.}, 3(4):510--530, 2009.
%\newblock MR2561668.
\bibitem{[16]Kumam2009newHybridIterativeMethod}
M.~Lashkarizadeh Bami and E.~Soori.
\newblock Strong convergence of a general implicit algorithm for variational
  inequality problems and equilibrium problems and a continuous representation
  of nonexpansive mappings.
\newblock {\em Bull. Iranian Math. Soc.}, 40(4):977--1001, 2014.
%\newblock MR3255412.
\bibitem{[28]Onjae-uea2012convergenceOfInterativeSequences}
N.~Onjae-uea, C.~Jaiboon, and P.~Kumam.
\newblock Convergence of iterative sequences for fixed points of an infinite
  family of nonexpansive mappings based on a hybrid steepest descent methods.
\newblock {\em J. Inequal. Appl.}, 101:22 pp., 2012.
%\newblock MR2931014.

\bibitem{[17]Peathanom2011hybridIterativeMethodGeneralizedEquilibrium}
S.~Peathanom and W.~Phuengrattana.
\newblock A hybrid method for generalized equilibrium, variational inequality
  and fixed point problems of finite family of nonexpansive mappings.
\newblock {\em Thai J. Math.}, 9(1):95--119, 2011.
%\newblock MR2833756.

\bibitem{[18]Piri2012generalIterativeMethodEquilibriumSystems}
H.~Piri.
\newblock A general iterative method for finding common solutions of system of
  equilibrium problems, system of variational inequalities and fixed point
  problems.
\newblock {\em Math. Comput. Modelling}, 55(3-4):1622--1638, 2012.
%\newblock MR2887545.

\bibitem{[20]Qin2008generalIterativeMethodForEquilibriumProblems}
X.~Qin, M.~Shang, and Y.~Su.
\newblock A general iterative method for equilibrium problems and fixed point
  problems in Hilbert spaces.
\newblock {\em Nonlinear Anal.}, 69(11):3897--3909, 2008.
%\newblock MR2463342.

\bibitem{[21]Qin2008generalIterativeMethodForEquilibriumProblems}
X.~Qin, M.~Shang, and Y.~Su.
\newblock Strong convergence of a general iterative algorithm for equilibrium
  problems and variational inequality problems.
\newblock {\em Math. Comput. Modelling}, 48(7--8):1033--1046, 2008.
%\newblock MR2458216.

\bibitem{[19]Qin2008strongConvergence}
X.~Qin, M.~Shang, and H.~Zhou.
\newblock Strong convergence of a general iterative method for variational
  inequality problems and fixed point problems in Hilbert spaces.
\newblock {\em Appl. Math. Comput.}, 200(1):242--253, 2008.
%\newblock MR2421640.
 \bibitem{sa} S. Saeidi, Comments on relaxed $(\gamma, r)$-cocoercive mappings. Int. J. Nonlinear Anal. Appl.
1 (2010) No.1, 54-57.
\bibitem{[24]Saewan2011shrinkingProjection}
S.~Saewan and P.~Kumam.
\newblock The shrinking projection method for solving generalized equilibrium
  problems and common fixed points for asymptotically quasi-$\phi$-nonexpansive
  mappings.
\newblock {\em Fixed Point Theory Appl.}, 9:25 pp., 2011.
%\newblock MR2820313.

\bibitem{[25]Shan2012iterativeMethodVectorEquilibrium}
S.~Shan and N.~Huang.
\newblock An iterative method for generelized mixed vector equlibrium problems
  and fixed point of nonexpansive mappings and variathional inequalities.
\newblock {\em Taiwanese Journal of Mathematics}, 16:1681--1705, 2012.

\bibitem{[26]Sun2014hybridMethods}
L.~Sun.
\newblock Hybrid methods for common solutions in Hilbert spaces with
  applications. (english summary).
\newblock {\em J. Inequal. Appl.}, 183(MR3239824):16 pp., 2014.

\bibitem{[27]Takahashi2008strongConvergenceTheorem}
S.~Takahashi and W.~Takahashi.
\newblock Strong convergence theorem for a generalized equilibrium problem and
  a nonexpansive mapping in a Hilbert space.
\newblock {\em Nonlinear Anal.}, 69(3):1025--1033, 2008.
%\newblock MR2428774.

\bibitem{[27b]Takahashi2013inverseStronglyMonotone}
W.~Takahashi.
\newblock Strong convergence theorems for maximal and inverse-strongly monotone
  mappings in Hilbert spaces and applications.
\newblock {\em J. Optim. Theory Appl.}, 157(3):781--802, 2013.
%\newblock MR3047030.

\bibitem{[29]Yao2010variationalInequalityProblems}
Y.~Yao, Y.~Liou, and S.~Kang.
\newblock Approach to common elements of variational inequality problems and
  fixed point problems via a relaxed extragradient method.
\newblock {\em Comput. Math. Appl.}, 59(11):3472--3480, 2010.
%\newblock MR2646318.

\bibitem{[30]Zhang2012stronglyMonotone}
M.~Zhang.
\newblock Iterative algorithms for common elements in fixed point sets and zero
  point sets with applications.
\newblock {\em Fixed Point Theory Appl.}, 21:14 pp., 2012.
%\newblock MR2899784.

\bibitem{[30]Zhang2012iterativeAlgorithmsFixedPointSets}
M.~Zhang.
\newblock Iterative algorithms for common elements in fixed point sets and zero
  point sets with applications.
\newblock {\em Comput. Math. Appl.}, 21:14 pp., 2012.
%\newblock MR2899784.

\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage{amssymb,amsmath,latexsym}

\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, graphicx, color}
\usepackage{marvosym}
\usepackage{paralist}
\usepackage{color}

\newcommand{\assign}{\mathrel{\mathop :}=}
\usepackage[varg]{pxfonts}

\usepackage{stmaryrd}

\usepackage{paralist}

\newtheorem{lem}{\textbf{Lemma}}[section]

\newtheorem{theorem}[lem]{\textbf{Theorem}}

\newtheorem{rem}[lem]{\textbf{Remark}}

\theoremstyle{definition}

\theoremstyle{definition}
\theoremstyle{remark}

\newcommand{\norm}[1]{\left\Vert#1\right\Vert}

\newcommand{\abs}[1]{\left\vert#1\right\vert}

   \thanks{\!\!\!\!\!\!\!\! \!\!${^{*}}$Corresponding author \\  2010 Mathematics Subject Classification: 47H09; 47H10.\\Keywords:  Inverse-strongly monotone mapping; Strongly monotone mapping; Relaxed $(u, v)$-cocoercive mapping. \\ E-mail addresses: agarwal@tamuk.edu, sori.e@lu.ac.ir,\\ donal.oregan@nuigalway.ie  }

\begin{document}
\title[A Simple proof  for  the algorithms ]{A Simple proof  for  the algorithms of relaxed $(u, v)$-cocoercive mappings and  $\alpha$-inverse strongly monotone mappings}
\author[ Agarwal,   Soori,   O'Regan ]{ Ravi P. Agarwal${^{1}}$, Ebrahim Soori${^{*,2}}$, Donal O'Regan${^{3}}$}


\date{}


%

\begin{abstract}
In this paper, a simple proof   is   presented    for the Convergence of the algorithms for the class of relaxed $(u, v)$-cocoercive mappings and $\alpha$-inverse strongly monotone mappings.    Based on $\alpha$-expansive maps, for example,
    a  simple  proof of the convergence of  the recent iterative algorithms  by relaxed $(u, v)$-cocoercive mappings   due to  Kumam-Jaiboon is provided.      Also    a simple  proof for the convergence of  the   iterative algorithms by inverse-strongly monotone mappings  due to Iiduka-Takahashi in a special case   is provided.   These results are an improvement as well as a refinement of previously known results.
\end{abstract}


\maketitle


\section{ Introduction}
In this paper,    some  results for the class of  relaxed $(u, v)$-cocoercive mappings and $\alpha$-inverse strongly monotone mappings (in a spacial case) are presented. There are many papers in the literature on  iterative algorithms which are used  for example in optimization problems, variational inequality problems, fixed point problems, equilibrium problems, Nash equilibrium problems, game theory, saddle point problems, minimization problem, feasibility problems, complementarity problems; see  \cite{[15]Kumam2009newHybridIterativeMethod} and the references therein.
In this paper,    a  simple proof   for the convergence of recent iterative algorithms is  presented which improve and refine the proof of  known results in the literature.

\section{Preliminaries}
In this paper,  it is assumed that   $C$ is  a nonempty closed convex subset  of a real Hilbert space $H$ with inner product $\langle \cdot, \cdot \rangle$ and norm $\norm{\cdot}$.    Recall the following well known  concepts:
\begin{enumerate}
   \item a mapping   $A : C \rightarrow H $   is said to be inverse-strongly monotone~\cite{[6]Jaiboon2010cocoerciveMappings}~\cite{[27b]Takahashi2013inverseStronglyMonotone}, if there exist $\alpha > 0$ such that
 \begin{equation*}
    \langle Ax - Ay, x - y\rangle \geq \alpha\|Ax - Ay\|^{2},
 \end{equation*}
   for  all  $x, y \in C$.
   \item  a mapping   $A : C \rightarrow H $   is said to be strongly monotone~\cite{[30]Zhang2012stronglyMonotone}~\cite[\S 1, p. 200]{BrowderPetryshyn1967inverseStronglyMonotone}, if there exists a constant $\alpha > 0$ such that
\begin{equation}\label{hjgfd}
      \langle Ax - Ay, x - y\rangle \geq \alpha\| x -  y\|^{2},
\end{equation}
\item  a mapping $B: C \rightarrow H $ is  said to be relaxed $(u, v)$-cocoercive~\cite{[6]Jaiboon2010cocoerciveMappings}, if there exist two constants $u, v > 0$ such that
 \begin{equation*}
    \langle Bx - By , x - y\rangle \geq (-u)\|Bx - By\|^{2}+v\|x - y\|^{2},
 \end{equation*}
   for  all  $x, y \in C$. For $u = 0$, $B$ is $v$-strongly monotone
	~\cite{[30]Zhang2012stronglyMonotone}.   Clearly, every  $v$-strongly monotone map is a relaxed $(u, v)$-cocoercive map.
\item for a map  $B: C \rightarrow H $ the classical variational problem is to find a $u\in C$ such that   $\langle Bu, v - u\rangle \geq 0,\,\,\forall v\in C$. We denote by  $V I(C,B)$
the set of solutions of the variational inequality problem,
\item let $C$ be a nonempty closed convex subset of a real Hilbert space $ H $.
Let $B$ be a  self  mapping on $C$.  Suppose that  there exists      a positive  integer $\alpha$ such that
 \begin{equation*}\label{golf2}
      \| Bx -By\| \geq \alpha\| x -  y\|.
  \end{equation*}
  for  all  $x, y \in C$, then $B$ is said to be $\alpha$-expansive.
\end{enumerate}

In this paper, using   relaxed $(u, v)$-cocoercive mappings,    a new  proof of some  recent iterative algorithms is presented. A similar comment applies for  inverse-strongly monotone mappings.

\section{\textbf{Relaxed $(u, v)$-cocoercive mappings}}

G. Cai and  S. Bu \cite{[1]CaiBu2011strongConvergence},  W. Chantarangs, C. Jaiboon, and P. Kumam \cite{[2]Chantaragsi2010coerciveMappings},   J.S. Jung \cite{[11]Jung2011compositeIterativeMethod},   P. Kumam and  C. Jaiboon \cite{[15]Kumam2009newHybridIterativeMethod},   X. Qin,  M. Shang and H. Zhou \cite{[19]Qin2008strongConvergence},    X. Qin, M. Shang and Y. Su \cite{[21]Qin2008generalIterativeMethodForEquilibriumProblems},  X. Qin, M. Shang and Y. Su \cite{[20]Qin2008generalIterativeMethodForEquilibriumProblems},
considered some iterative algorithms  for finding a common element of the set of
fixed points of nonexpansive mappings and the set of  solutions of a variational
inequality $V I(C,A)$, where $A$ is a relaxed $(u, v)$-cocoercive  mapping
of $C$ into $H$.

\begin{lem}\label{sgjm}
     Let $A$ be    a relaxed $(m, v)$-cocoercive mapping  and $\epsilon$-Lipschitz continuous such that $v-m \epsilon^{2}>0$ and $V I(C,A)\neq \emptyset$. Then  $A$  is    a   $(v-m \epsilon^{2})$-expansive mapping  and   $V I(C,A)$ is a singleton.
\end{lem}
\begin{proof}
To see that  $V I(C,A)$ is a singleton, one can see  \cite[Proposition 2]{sa}.
Next, since $A$ is $(m, v)$-cocoercive and  $\epsilon$-Lipschitz continuous, for each $x,y \in C$, it is concluded that
\begin{align*}\label{jhgjgl}
  \langle Ax-Ay, x-y \rangle& \geq(-m)\| Ax-Ay\|^{2}+v\|x-y \|^{2} \nonumber \\ \nonumber &\geq (-m \epsilon^{2})\|x-y \|^{2} +v\|x-y \|^{2}\\  &= (v-m \epsilon^{2})\|x-y \|^{2}\geq 0,
\end{align*}
   hence, we have
 \begin{equation*}
   \| Ax-Ay\|\geq  (v-m \epsilon^{2})\|x-y \|,
 \end{equation*}
so $A$ is   $(v-m \epsilon^{2})$-expansive.
\end{proof}

To see an example,      Theorem 3.1 from P. Kumam and C. Jaiboon \cite{[15]Kumam2009newHybridIterativeMethod} is considered. To solve the mixed equilibrium problem for an equilibrium function $\Theta:E\times E\longrightarrow \mathbb{R}$, assume ~\cite[\S 2, p. 512]{[15]Kumam2009newHybridIterativeMethod} that $\Theta$ satisfies the following conditions:

\begin{compactenum}[(H1)]
\item $\Theta$ is monotone, {\em i.e.}, $\Theta(x,y) + \Theta(y,x) \leq 0, \forall x,y\in E$.
\item For each fixed $y\in E, x\mapsto \Theta(x,y)$ is convex and upper semicontinuous.
\item For each $x\in E, x\mapsto \Theta(x,y)$ is convex.
\end{compactenum}

\begin{theorem}\label{thm:3.1}{ (\rm {\em i.e.}, Theorem 3.1, from~\cite[\S 3, p.515]{[15]Kumam2009newHybridIterativeMethod})}
Let $E$ be a nonempty closed convex subset of a real Hilbert space $H$ and let $\varphi$ be a lower semicontinuous  and convex functional from $E$ to $\mathbb{R}$.   Let $\Theta$ be a bifunction from $E\times E$ to $\mathbb{R}$ satisfying $(H1)-(H3)$, let
$\left\{T_n\right\}$ be an infinite family of nonexpansive mappings of $E$ into itself and let $B$ be a $\xi$-Lipschitz continuous and relaxed $(m,v)$-cocoercive map of $E$ into $H$ such that
\[
\Gamma \assign \mathop{\bigcap}\limits_{n=1}^{\infty}F\left(T_n\right)\cap \Omega\cap VI(E,B)\neq \emptyset;
\]
here $\Omega$ is the set of solutions of the mixed variational probem (i.e. the set of $x$'s in $E$ with $\Theta(x,y)+\varphi(y)-\varphi(x) \geq 0,\,\,\forall y \in E$) and $F(T_n)$ is the set of fixed points of $T_n$).
Let $\mu > 0, \gamma > 0, r > 0$, be three constants.   Let $f$ be a contraction of $E$ onto itself with $\alpha\in (0,1)$ and let $A$ be a strongly positive linear bounded operator on $H$ with coefficient $\bar{\gamma} > 0$ and $0 < \gamma < \frac{(1+\mu)\bar{\gamma}}{\alpha}$.   For $x_1\in H$ arbitrarily and fixed $u\in H$, suppose $\left\{x_n\right\}, \left\{y_n\right\}$ and $\left\{z_n\right\}$ are generated iteratively by
\begin{equation*}
\begin{cases}
\Theta(z_n,x) + \varphi(x) - \varphi(z_n) + \frac{1}{r}\langle K'(z_{n})- K'(x_{n}), \eta(x,z_{n} )\rangle \geq 0,\\
y_n = \alpha_n z_n + (1 - \alpha_n)W_nP_E\left(z_n - \lambda_n B z_n\right),\\
x_{n+1} = \epsilon_n\left(u + \gamma f(W_nx_n)\right) + \beta_n x_n + \left((1- \beta_n)I - \epsilon_n(I + \mu A)\right)W_nP_{E}(y_{n}-\tau_{n}By_{n}), &\text{\quad}\\
\end{cases}
\end{equation*}
for all $n\in \mathbb{N}$ and   $ x\in E$, where $W_n$ is the $W$-mapping defined by (2.3) in ~\cite[\S 2, p. 514]{[15]Kumam2009newHybridIterativeMethod} and $\{\epsilon_n\},\{\alpha_n\}$ and $\{\beta_n\}$ are three sequences in (0,1).  Assume the following conditions are satisfied:
\begin{compactenum}[(C1)]
\item $\eta:E\times E \longrightarrow H$ is Lipschitz continuous with constant $\lambda > 0$ such that:
 \begin{compactenum}[(a)]
  \item $\eta(x,y) + \eta(y,x) = 0, \forall x,y\in E$,
  \item $\eta(\cdot,\cdot)$ is affine in the first variable,
  \item for each fixed $y\in E, x\mapsto \eta(y,x)$ is sequentially continuous from the weak topology to the weak topology;
  \end{compactenum}
\item $K: E\longrightarrow \mathbb{R}$ is $\eta$-strongly convex with constant $\sigma > 0$ and its derivative $K'$ is not only sequentially continuous from the weak topology to the strong topology but also Lipschitz continuous with constant $\nu > 0$ such that $\sigma > \lambda\nu$;
\item for each $x\in E$, there exists a bounded subset $D_x \subset E$ and $z_x\in E$ such that, for any $y\in E\setminus D_x$,
\[
\Theta(y, z_x) + \varphi(z_x) - \varphi(y) + \frac{1}{r}\langle K'(y) - K'(x), \eta(z_x,y)\rangle < 0;
\]
\item $\mbox{lim}_{\eta\rightarrow \infty}\alpha_n = 0, \mbox{lim}_{\eta\rightarrow \infty}\epsilon_n = 0$ and $\sum_{n=1}^{\infty}\epsilon_n = \infty$;
\item $0 < \mbox{lim inf}_{\eta\rightarrow \infty}\beta_n \leq \mbox{lim sup}_{\eta\rightarrow \infty}\beta_n < 1$;
\item $\mbox{lim}_{\eta\rightarrow \infty}\abs{\lambda_{n+1}-\lambda_{n}} = \mbox{lim}_{\eta\rightarrow \infty}\abs{\tau_{n+1}-\tau_{n}} = 0$;
\item $\left\{\tau_n\right\},\left\{\lambda_n\right\}\subset [a,b]\ \mbox{for some}\ a,b\ \mbox{with}\ 0\leq a\leq b\leq \frac{2\left(\upsilon-m\xi^2\right)}{\xi^2}.$
\end{compactenum}

Then $\left\{x_n\right\}$ and $\left\{z_n\right\}$ converge strongly to $z\in \Gamma \assign \bigcap_{n=1}^{\infty}F(T_n)\cap \Omega\cap VI(E,B)$, provided that $S_r$ (here $S_r$ is given in  ~\cite[\S 2, p. 513]{[15]Kumam2009newHybridIterativeMethod}) is firmly nonexpansive, which solves the following optimization problem:
\[
OP: \mathop{\mbox{min}}\limits_{x\in\Gamma}\frac{\mu}{2}\langle Ax,x\rangle +\frac{1}{2}\norm{x - u}^2 - h(x);
\]
here $h$ is a potential function for $\gamma\,f$.
\end{theorem}
Next, in the following remark, a simple proof for some similar results is presented:
  \begin{rem}   \textbf{A simple Proof:}\label{bhsskkl54ww}
(i). Consider Theorem~\ref{thm:3.1}  and the  $\xi$-Lipschitz continuous and   relaxed $(m, v)$-cocoercive mapping   $B$ in  Theorem ~\ref{thm:3.1}. From
condition (C7) we may assume that $(v-m\xi^{2})>0$, and hence from Lemma \ref{sgjm}, $B$  is  $(v-m \epsilon^{2})$-expansive,
i.e,
 \begin{equation}\label{djjkmk}
      \| Bx -By\| \geq (v-m\xi^{2})\|x - y\|,
  \end{equation}
and $VI(E, B) $ is   singleton i.e.  there exists an element $p \in E$ such that $VI(E, B)=\{p\} $, hence  $\Gamma=\{p\} $ in Theorem ~\ref{thm:3.1}.
  The authors    prove (see  (3.25) in ~\cite[ p. 521]{[15]Kumam2009newHybridIterativeMethod}) that
  \begin{equation}\label{golf2ww}
    \displaystyle \lim_{n}\|Bz_{n}-Bp\|=0.
  \end{equation}
 Now, put  $x=z_{n}$  and $y=p$ in  \eqref{djjkmk},  and  from \eqref{djjkmk} and \eqref{golf2ww},  we have
   \begin{equation*}
     \displaystyle \lim_{n}\|z_{n}-p\|=0.
   \end{equation*}
Hence,    $z_{n} \rightarrow p$.  As a result one of the main claims of  Theorem  ~\ref{thm:3.1}  is established (note $ \Gamma= VI(E, B)=\{p\} $).
Note the proof in Theorem ~\ref{thm:3.1} can  be simplified further by using this remark (for example Step 5 in ~\cite[ p. 526-528]{[15]Kumam2009newHybridIterativeMethod} is not needed since one can deduce it from
 (3.33)  in  ~\cite[ p. 524]{[15]Kumam2009newHybridIterativeMethod} and  the fact that  $z_{n} \rightarrow p$).

(ii). A similar remark applies to the main results in  \cite{[1]CaiBu2011strongConvergence,
[2]Chantaragsi2010coerciveMappings,  [19]Qin2008strongConvergence,[20]Qin2008generalIterativeMethodForEquilibriumProblems,[21]Qin2008generalIterativeMethodForEquilibriumProblems}.

 \end{rem}

\section{\textbf{Inverse strongly monotone mappings}}
J. Chen, L. Zhang, T. Fan \cite{[3]Chen2007nonexpansiveMappings}, S. Takahashi and W. Takahashi \cite{[27]Takahashi2008strongConvergenceTheorem}, H. Iiduka and W.Takahashi \cite{[4]IidukaTakahashi2005strongConvergenceTheorems},  K. R. Kazmi,  Rehan Ali and Mohd Furkan \cite{[14]Kazmi2018hybridIterativeMethod},  X. Qin, M. Shang and Y. Su \cite{[21]Qin2008generalIterativeMethodForEquilibriumProblems},  M. Zhang \cite{[30]Zhang2012iterativeAlgorithmsFixedPointSets},  S. Shan and N. Huang \cite{[25]Shan2012iterativeMethodVectorEquilibrium}, T. Jitpeera and  P. Kumam \cite{[7]Jitpeera2011fixedPointProblems},  S. Peathanom and W. Phuengrattana\cite{[17]Peathanom2011hybridIterativeMethodGeneralizedEquilibrium}, Piri \cite{[18]Piri2012generalIterativeMethodEquilibriumSystems},  X. Qin, M. Shang and Y. Su \cite{[20]Qin2008generalIterativeMethodForEquilibriumProblems} and  M. Lashkarizadeh Bami and E. Soori \cite{[16]Kumam2009newHybridIterativeMethod}
considered some iterative methods for finding a common element of a set of
fixed points of nonexpansive mapping and the set of  solutions of a variational
inequality $V I(C,A)$, where $A$ is an $\alpha$-inverse strongly monotone mapping
of $C$ into $H$. In this section, a  spacial case similar to cocoercive mappings is studied.
\begin{lem}\label{qwovnrt}
   Let $A$ be     an  $\alpha$-inverse strongly monotone mapping and  $V I(C,A)\neq \emptyset$.  Suppose that  $A$  is  also an  $\gamma$-expansive mapping. Then $V I(C,A)$ is a singleton.
\end{lem}
\begin{proof}
  Since $A$ is an  $\alpha$-inverse strongly monotone mapping, it is implied that
  \begin{equation}\label{gdwzl2}
        \langle Ax - Ay, x - y\rangle \geq \alpha\|Ax - Ay\|^{2}\geq 0,
  \end{equation}
(so $A$ is monotone).
  Since $A$ is $\gamma$-expansive, it is concluded  that
   \begin{equation}\label{kkfc2}
    \|Ax - Ay\| \geq \gamma \|x-y\|.
  \end{equation}
   Therefore,   $A$  is one to one, because if $Ax=Ay$, then  from \eqref{kkfc2}, $\|x-y\|=0$, and hence $x=y$.
Let $x_{1}, x_{2} \in V I(C,A)$. Then
\begin{equation}\label{lbjeo2}
  \langle Ax_{1}, y-x_{1}  \rangle \geq 0,
\end{equation}
for each $y \in C$, and
\begin{equation}\label{bjeo2}
  \langle Ax_{2}, y-x_{2}  \rangle \geq 0,
\end{equation}
for each $y \in C$. Substitute $x_{1}$ in  \eqref{bjeo2} and  $x_{2}$ in  \eqref{lbjeo2}, then     $\langle Ax_{1}, x_{2}-x_{1}  \rangle \geq 0$ and
 $ \langle Ax_{2}, x_{1}-x_{2}  \rangle \geq 0$. Adding them, it is implied that
\begin{equation}\label{qwpld2}
  \langle Ax_{2}-Ax_{1}, x_{2}-x_{1}  \rangle \leq 0.
\end{equation}
Since  $A$ is monotone,
 then   $  \langle Ax_{2}-Ax_{1}, x_{2}-x_{1}  \rangle \geq 0$,  and hence from \eqref{qwpld2}, it is concluded  that
\begin{equation}\label{ovl}
    \langle Ax_{2}-Ax_{1}, x_{2}-x_{1}  \rangle = 0.
\end{equation}
Then from \eqref{gdwzl2},  $Ax_{2}=Ax_{1}$ and  since $A$ is one to one, it is gotten that $x_{2}=x_{1}$. Then $V I(C,A)$ is   singleton.
\end{proof}

\begin{rem} Note in Lemma~\ref{qwovnrt} we can replace (a): $A$ being a   $\gamma$-expansive mapping  with $A$ being one to one, and (b): $A$ being an
  $\alpha$-inverse strongly monotone  with $A$ being monotone.
\end{rem}
	
\begin{rem} \textbf{A simple proof in a spacial case:}\label{bhsskkl}
(i).  Consider   Theorem 3.1 in  \cite[\S 3, p. 343]{[4]IidukaTakahashi2005strongConvergenceTheorems}; note   $A$ is an  $\alpha$-inverse strongly monotone mapping there. If we consider    an extra condition  of $\gamma$-expansiveness  (or $A$ being one to one)  in Theorem 3.1   in  \cite{[4]IidukaTakahashi2005strongConvergenceTheorems}
 then from Lemma \ref{qwovnrt}, we have  that $V I(C,A)$ is a singleton and hence   in  Theorem 3.1 in  \cite[\S 3, p. 343]{[4]IidukaTakahashi2005strongConvergenceTheorems} $F(S)\cap V I(C,A)$   is a  singleton, i.e,   $F(S)\cap V I(C,A)=V I(C,A)=\{u\}$ for an element $u\in C$.  The authors   prove  (see  line 8  from below in
 \cite[ p. 345]{[4]IidukaTakahashi2005strongConvergenceTheorems}) that
  \begin{equation}\label{golf}
    \displaystyle \lim_{n}\|Ax_{n}-Au\|=0.
  \end{equation}
  Now, put  $x=x_{n}$  and $y=u$   and then from \eqref{golf} and \eqref{golf2},  we have
   \begin{equation*}
     \displaystyle \lim_{n}\|x_{n}-u\|=0.
   \end{equation*}
Hence,    $x_{n} \rightarrow u$.   Therefore   we have     the main claim of  Theorem 3.1 (note  $ F(S)\cap V I(C,A)= VI(E, B)=\{u\} $). As a result in this situation  one
can remove everything in the proof after line 8 from below    in  \cite[ p. 345]{[4]IidukaTakahashi2005strongConvergenceTheorems}.

(ii). A similar comment applies to the main results in   \cite{[3]Chen2007nonexpansiveMappings, [4]IidukaTakahashi2005strongConvergenceTheorems,[7]Jitpeera2011fixedPointProblems,[8]Jitpeera2011equilibriumProblems,[9]Jitpeera2011hybridAlgorithms,[10]Jitpeera2011shrinkingProjection,[11]Jung2011compositeIterativeMethod,[12]Kangtunyakarn2010iterativeMethods,[13]Kazmi2012extragradientMethod,[14]Kazmi2018hybridIterativeMethod,[16]Kumam2009newHybridIterativeMethod,[17]Peathanom2011hybridIterativeMethodGeneralizedEquilibrium,[18]Piri2012generalIterativeMethodEquilibriumSystems,[20]Qin2008generalIterativeMethodForEquilibriumProblems,[21]Qin2008generalIterativeMethodForEquilibriumProblems,[24]Saewan2011shrinkingProjection,[25]Shan2012iterativeMethodVectorEquilibrium,[26]Sun2014hybridMethods,[27]Takahashi2008strongConvergenceTheorem,[28]Onjae-uea2012convergenceOfInterativeSequences,[29]Yao2010variationalInequalityProblems,[30]Zhang2012iterativeAlgorithmsFixedPointSets} is similar   to \cite{[4]IidukaTakahashi2005strongConvergenceTheorems}.

 \end{rem}
\begin{rem}
Note that the proof of  these algorithms for inverse  strongly monotone mappings  can not be     simplified like in  the proof of algorithms for cocoercive mappings in general. Indeed in   Remark \ref{bhsskkl},  a simple proof is just introduced for a special case of these algorithms.
\end{rem}

\section{Discussion}
In this paper, a simple  proof   for the convergence of  the algorithms  of relaxed $(u, v)$-cocoercive
mappings and $\alpha$-inverse strongly monotone mappings are presented.
Indeed, a refinement of the proofs of some well known results
are presented.
\section{Conclusion}
In this paper, a refinement  of the proofs of some well known
results are given. Indeed, the proofs of these results are made shorter
than the original ones for the class of relaxed $(u, v)$-cocoercive mappings
and $\alpha$-inverse strongly monotone mappings.
\section{Acknowledgements}
The second author is grateful to the University of Lorestan for their
support.
\section{Funding}
Not applicable
\section{Abbreviations}
Not applicable
\section{Availability of data and material}
Please contact author for data requests.
\section{Competing interests}
The authors declare  that they  have no competing interests.
\section{ Authors’ contributions}
All authors contributed equally to the manuscript, read and approved
the final manuscript.

\section{ Author details}
$^{1}$Department of
Mathematics, Texas A  \& M University Kingsville, Kingsville, USA. $^{2}$Department of Mathematics, Lorestan University,
P.O. Box 465, Khoramabad, Lorestan, Iran.  $^{3}$School of Mathematics, Statistics, National University of Ireland,
Galway, Ireland.

\end{document}